OpenStudy (anonymous):

Given the equation f(x)=1/(x-2) -3, I can't find any critical values so is this a discontinuous equation?!

OpenStudy (aum):

\[ f(x) = \frac{1}{(x-2)} - 3 \\ f'(x) = -\frac{1}{(x-2)^2} \\ \text{There is no solution to }f'(x) = 0.\\ f'(x) \text { is undefined at x = 2.} \\ \text{However, x = 2 is NOT a critical point because f(x) does not exist at x = 2 either.}\\ \text{No critical points for this function.} \]